OpenStudy (anonymous):

A 400.0g metal bar requires 6.446 kJ of energy to change its temperature from 18.0ºC to 95.0ºC. What is the specific heat of the metal?

OpenStudy (anonymous):

use Q=m*c*(t2-t1)

OpenStudy (anonymous):

can u do now

OpenStudy (anonymous):

@kphilxo Q=6.446 , m=0.4 kg , t =95-18=77 can u do nw

OpenStudy (anonymous):

yes thank you